A rock with a mass of 6.0 kg is being whirled in a horizontal circle on the end of a rope 1.6m long. If the speed of the rock is 2.7m/s , what is the magnitude of the centripetal acceleration?
1
Expert's answer
2011-12-13T11:19:28-0500
The liniar velocity of a point at the tub is 2.7 m/s, so centripetal acceleration is
Ac = V²/R = 2.7²/1.6 = 4.55625 m/s².

The magnitude of the centripetal acceleration is called the centrifugal force and it is
F = M*Ac = 6.0*4.55625 = 27.3375 N.